OpenStudy (anonymous):

write an equation for the line with the give condition that contains the point (-1,2) Parallel to the line y=x-6

OpenStudy (anonymous):

pralledl to line y=x-6 means slope are equal........so slope equals to 1 using (y-y1)=m(x-x1) y-2=1(x+1) =y=x+3

OpenStudy (anonymous):

its y=x+3

OpenStudy (anonymous):

thankyou! what would be the perpendicular to the line y=-x?

OpenStudy (anonymous):

slope would be -1/slope of the line i.e. -1

OpenStudy (anonymous):

isn't it suppose to be an equation?

OpenStudy (anonymous):

u put the value of slope as -1 in y-y1 =m(x-x1) but same co-ordinates value